OpenStudy (anonymous):

Help! i I can not do inverse matrix by hand! can anyone give me tips? thanks!

OpenStudy (anonymous):

What are the dimensions of the matrix?

OpenStudy (anonymous):

3x3 - 1st row 4 2 2 2nd row 4 2 0 3rd row 5 0 5

OpenStudy (anonymous):

\[\left[\begin{matrix}4 & 2 & 2 \\4 & 2 & 0\\ 5 & 0 & 5\end{matrix}\right]\] Do you know how to take the determinant of a two-by-two matrix?

OpenStudy (anonymous):

I guess not.

OpenStudy (anonymous):

I was asking to check. I can help you. If you have a two-by-two:\[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] The determinant of the matrix is:\[a*d-b*c\]

OpenStudy (anonymous):

I can do a 2x2 determinent. Actually not sure how to do an inverse on a 2x2 either :-)

OpenStudy (anonymous):

I realize now that I was about to get you the determinant of the three-by-three and not the inverse... It's difficult to inverse high-order matrices, do you know how to row-reduce?

OpenStudy (anonymous):

I know you have to find the identity using gauss-jordan, but I don't know how they keep finding the inverse from that.

OpenStudy (anonymous):

I have no idea what gauss-jordan is.\[\left[\begin{matrix}-1/2 & 1/2 & 1/5 \\1&-1/2&-2/5\\ 1/2 & -1/2&0\end{matrix}\right]\] I found that using wolfram alpha.

OpenStudy (anonymous):

THANKS FOR YOUR HELP. SORRY I DIDN'T REPLY SOONER, AT WORK AND PHONE WENT CRAZY.

OpenStudy (anonymous):

Don't worry about it :)